Question 15: In a Cournot market, how does a firms output change as we add more firms to the market?
Answer #1

If the market demand function be: P = a - bQ where Q = q1 + q2 + q3 +......qn where n: number of firms, and

Marginal cost = c (constant), then

Equilibrium quantity (Q) = [n x (a - c)] / [b x (n + 1)]

Since a, b, c > 0, as n increases, (n + 1) increases by a higher amount and so, denominator of above expression increases more than proportionately than the numerator. Therefore, Q decreases as n increases.